On a 100-mile trip,a family drove for 1 hour at 50 mph and then the rest at 60 mph. What was the average speed of the trip?
Answer by stanbon(75887)

On a 100-mile trip,a family drove for 1 hour at 50 mph and then the rest at 60 mph. What was the average speed of the trip?
--------------
Average speed = (total distance)/(total time)
------------------
1 hr at 50 mph = 50 miles
---
That leaves 50 miles to go at 60 mph.
---
time = d/r = 50/60 = 5/6 hr
--------------------------------
Equation:
Average speed = 100 miles/ (1hr + 5/6 hr) = 100/(11/6) = 600/11 = 54.5455 mph
=============================================================================
Cheers,
Stan H.